OpenStudy (anonymous):

Find the y-intercept of the line whose equation is 3x - 4y = -2. im lost

OpenStudy (anonymous):

To find the y-intercept: Rearrange the equation to solve for y, then set x to 0. 3x - 4y - 12 = 0 4y = 3x - 12 y = (¾)x - 3 y = (¾)0 - 3 = -3